# Analyzing the numerical correctness of branch-and-bound decisions for mixed-integer programming

### v1

@@ 1 TL 1 1 0 0 0 0 0
tkatTV5 2.8117644.225 2.8117644.225 28117644.225
Table 5: Comparison of the results of floating-point SCIP and rational solutions. The exact results were generated with Exact-SCIP (hash: c7bd4be7b9) and a time limit of 18800 seconds
As expected, on the numerical more challenging NumDiff test, more wrong decisions are made. First, let us focus on the instances with dual fails. Besides the instances alu16_8, tkatTV5, and tkat3TV, all instances show at most two leaves with strong bound or gap errors. On the instances, tkat3*, ns1629327 and ns1859355, more weak and strong bound or gap errors appear, and alu10_9 produces 46 infeasibility errors. Nevertheless, these all makeup only a small fraction of the total number of evaluated nodes.
On the alu*-instances (except for alu16_1), and on the instances dfn6_load, neos-1053591, neos-1062641, neos-1603965, ns1866531 the solutions generated by SCIP are slightly infeasible and can not be converted to exact solutions. Especially the alu instances are numerical difficult: Either the instance is infeasible or there is a huge gap between the floating-point solution and the solution of Exact-SCIP as can be seen in Table 5.
One natural attempt to reduce such solution errors is to call SCIP with a tightened feasibility tolerance. Per default, the feasibility tolerance is $10^{-6}$. We repeated our experiments but tightened the tolerance to $10^{-9}$. The results show a significant reduction in solution errors. On the instances alu10_1, alu10_8, alu10_9, bernd2, aim-200 and dfn6_load, SCIP finds an exactly integer-feasible solution, or they can be converted to such, or SCIP times out, but on the instances 16_7 or ns1859355, the floating-point solver still produces slightly infeasible solutions. On the FPEasy instance dano3_4 also all gap errors were removed, successfully closing the dual gap on this instance.
However, we also observe that the number of explored nodes is significantly increased. This does not only lead to increased solving time, but can also create more numerically challenging leaves, hence potentially increasing the absolute number of incorrect decisions. An example of this behavior is the instance blend2 where the result of SCIP can no longer be verified.
@@ 3.3 Analysis of different post-processing techniques
Finally, let us have a look at the effectiveness and success rate of the different post-processing techniques described in Section 2.3. Table 6 aggregates over all leaves of a complete run over each test set (column “leaves”) and categorizes them based on the techniques that were successful for verification. The column “floating-point” lists the percentage of leaves that are verified automatically or through the use of safe bounding. The columns “Reconstruct”, “Factorize”, and “Exact” likewise represent the percentage for the corresponding methods. The row “unb” reports the numbers for all instances that contain variables with unbounded domains.
@@
Surprisingly, 93.76% (FPEasy) to 99.69% (NumDiff) of the leaf decisions can be verified using floating-point arithmetic. Thus, the presented methods can be implemented very efficiently. In relative terms, the percentage of erroneous leaf decisions is small on all test sets.
Safe bounding requires bounded variables and can therefore more reliably be applied to problems with bounded variables. As a result, unbounded problems, listed in Table 6 in row “unb”, have a significantly higher percentage of expensive Exact-SoPlex calls. On these unbounded problems, the success rate of the floating-point arithmetic drops significantly to 65.18% on FPEasy and 94.89% on NumDiff. This leads to an increase in more expensive techniques such as the Exact-SoPlex calls which significantly increase from 2.85% to 29.89% on FPEasy and from 0.06% to 3.86% on NumDiff.

### v3

@@ 1 TL 1 1 0 0 0 0 0
tkatTV5 28117644.225 28117644.225 28117644.225
Table 5: Comparison of the results of floating-point SCIP and rational solutions. The exact results were generated with Exact-SCIP (hash: c7bd4be7b9) and a time limit of 18800 seconds.
As expected, on the numerically more challenging NumDiff test set, more wrong decisions are made. First, let us focus on the instances with dual fails. Besides the instances alu16_8, tkatTV5, and tkat3TV, all instances show at most two leaves with strong bound or gap errors. On the instances, tkat3*, ns1629327 and ns1859355, more weak and strong bound or gap errors appear, and alu10_9 produces 46 infeasibility errors. Nevertheless, these all make up only a small fraction of the total number of evaluated nodes.
On the alu*-instances (except for alu16_1), as well as on the instances dfn6_load, neos-1053591, neos-1062641, neos-1603965, ns1866531 the solutions generated by SCIP are slightly infeasible and can not be converted to exact solutions. Especially the alu instances are numerical difficult: Either the instance is infeasible or there is a huge gap between the floating-point solution and the solution of Exact-SCIP as can be seen in Table 5.
One natural attempt to reduce such solution errors is to call SCIP with a tightened primal feasibility tolerance. Per default, the feasibility tolerance is $10^{-6}$. We repeated our experiments but tightened the tolerance parameter “numerics/feastol” to $10^{-9}$. The results show a significant reduction in solution errors. On the instances alu10_1, alu10_8, alu10_9, bernd2, aim-200 and dfn6_load, SCIP finds an exactly integer-feasible solution, or they can be converted to such, or SCIP times out, but on the instances alu_16_7 or ns1859355, the floating-point solver still produces slightly infeasible solutions. On the FPEasy instance dano3_4 also all gap errors were removed, successfully closing the dual gap on this instance.
However, we also observe that the number of explored nodes is significantly increased. This does not only lead to increased solving time, but can also create more numerically challenging leaves, hence potentially increasing the absolute number of incorrect decisions. An example of this behavior is the instance blend2, for which the result of SCIP can no longer be verified after tightening the tolerance.
@@ 3.3 Analysis of Different Post-processing Techniques
Finally, let us have a look at the effectiveness and success rate of the different post-processing techniques described in Section 2.3. Table 6 reports aggregate results over all leaves of a complete run over each test set (column “leaves”) and categorizes them based on the techniques that were successful for verification. The column “floating-point” lists the percentage of leaves that are verified automatically or through the use of safe bounding. The columns “Reconstruct”, “Factorize”, and “Exact” likewise represent the percentage for the corresponding methods. The row “unb” reports the numbers for all instances that contain variables with unbounded domains.
@@
Surprisingly, 93.76% (FPEasy) to 99.69% (NumDiff) of the leaf decisions can be verified using floating-point arithmetic. Hence, the most successfull methods are also the ones that can be implemented most efficiently. In relative terms, the percentage of erroneous leaf decisions is small on all test sets.
Safe bounding requires bounded variables and can therefore more reliably be applied to problems with bounded variables. As a result, unbounded problems, listed in Table 6 in row “unb”, have a significantly higher percentage of expensive Exact-SoPlex calls. On these unbounded problems, the success rate of the “floating-point” techniques arithmetic drops significantly to 65.18% on FPEasy and 94.89% on NumDiff. This leads to an increase in more expensive techniques; most notably, the number of Exact-SoPlex calls significantly increases from 2.85% to 29.89% on FPEasy and from 0.06% to 3.86% on NumDiff.